Somethingserious15:

A telephone soliciting company obtains an average of five orders per 1000 solicitations. If the company reaches 250 potential customers, find the probability of obtaining at least two orders.

itsme:

any idea

SemiDefinite:

this is a single entry in a ClassWiz Binomial CD - p = 5/1000 - N = 250 - P(X >=2) = 1 - P(X<2) = 1 - P(X=1)